# Multiview Random Vector Functional Link Network for Predicting DNA-Binding Proteins

### v1

@@ A.1 Generalization Capability Analysis
If we multiply by $y_{i}$ on both sides of the constraints of Eq. (A.1), it results in
@@
[…]
Eqs. (38) and (39) are used in the theorem 2.
To start, we define the Rademacher complexity (Bartlett & Mendelson, 2002) as follows:
@@
For the set of samples, $S=\{x_{1},\ldots,x_{n}\}$, consisting of $n$ independent samples from the distribution $\mathcal{D}$ and the function set $\mathscr{G}$ on $S$, we define the empirical Rademacher complexity on $\mathscr{G}$ as:
@@
where $\sigma=(\sigma_{1},\ldots,\sigma_{n})$ are independently uniform valued $\{+1,-1\}$ (Rademacher) random variables. The Rademacher complexity of $\mathscr{G}$ is given by:
@@
Lemma 1.
Choose $\theta$ from the interval $(0,1)$ and consider $\mathscr{G}$ as a class of functions mapping from an input space $S$ to $[0,1]$. Suppose $\{x_{i}\}_{i=1}^{n}$ are drawn independently from a probability distribution $\mathcal{D}$. Then, with a probability of at least $1-\theta$ over random samples of size $n$, every $g\in\mathscr{G}$ satisfies:
@@
Lemma 3.
Consider $\mathscr{A}$ as a Lipschitz function with a Lipschitz constant $\mathcal{L}$, mapping the real numbers to the real numbers, satisfying $\mathscr{A}(0)=0$. The Rademacher complexity of the class $\mathscr{A}\circ\mathscr{G}$ is:
@@
We define the difference between the final predictors of the two views as $g_{m}(x)=\lvert Z_{1}^{(x)}\beta_{1}-Z_{2}^{(x)}\beta_{2}\rvert$. Thus, the true expectation bound of $g_{m}(x)$ can be derived using the following theorem.
Theorem 1.
Given $N\in\mathbb{R}^{+}$, $\theta\in(0,1)$, and a training set $T=\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn independently and identically from probability distribution $\mathcal{D}$, where $y_{i}\in\{-1,+1\}$ and $x_{i}=(x_{i}^{A};x_{i}^{B})$. Define the function class $\mathscr{G}_{N}=\{g|g:x\rightarrow\lvert Z^{(x)}\beta\rvert,\|\beta\|\leq N\}$ and $\hat{\mathscr{G}}_{N}=\{\hat{g}|\hat{g}:x\rightarrow Z^{(x)}\beta,\|\beta\|\leq N\}$, where $\beta=(\beta_{1};\beta_{2})$, $Z^{(x)}=\left(Z_{1}^{(x)};-Z_{2}^{(x)}\right)=\left([x^{A}\hskip 5.69046pt\phi(x^{A}W^{A}+b^{A})];-[x^{B}\hskip 5.69046pt\phi(x^{B}W^{B}+b^{B})]\right)$ and $g_{m}(x)=\lvert Z_{1}^{(x)}\beta_{1}-Z_{2}^{(x)}\beta_{2}\rvert=\lvert Z^{(x)}\beta\rvert\in\mathscr{G}_{N}$. Then, with a probability of at least $1-\theta$ over $T$, every $g_{m}(x)\in\mathscr{G}_{N}$ satisfies
@@
Let’s consider a loss function $\Omega:\mathbb{R}\rightarrow[0,1]$ defined as:
$\Omega=\left\{\begin{array}[]{lll}{-\frac{x}{N\mathcal{K}_{m}}},&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ -N\mathcal{K}_{m}\leq x<0,\vspace{3mm}\\ {\frac{x}{N\mathcal{K}_{m}}},&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ 0\leq x\leq N\mathcal{K}_{m},\\ {1},&\text{otherwise}.\end{array}\right.$ (50)
@@
In that case, $g_{m}(x)$ ranges from 00 to $N\mathcal{K}_{m}$, while $\hat{g}_{m}(x)$ ranges from $-N\mathcal{K}_{m}$ to $N\mathcal{K}_{m}$. According to Lemma 1 and with $\Omega(\hat{g}_{m}(x))$ ranging from 00 to $1$, the following inequality holds with a probability of at least $1-\theta$ over $T$:
@@
Given that $\Omega(x)$ is a Lipschitz function with a constant of $\frac{1}{N\mathcal{K}_{m}}$, passes through the origin, and is uniformly bounded, we can assert the following inequality based on Lemmas 2 and 3:
@@
From Eqs. (A.1) and (57), we can conclude the theorem. ∎
Consistency plays a crucial role in MVL, aiming to reduce the discrepancy in predictions across different perspectives or views. A lower consistency error typically results in improved generalization performance. We represent $g_{m}(x)$ as the consistency error function between the two views. Next, by applying the empirical Rademacher complexity of the function classes $\hat{\mathscr{G}}_{m}$ along with the empirical expectation of $g_{m}$, we obtain a margin-based estimate for the consistency error expectation. MvRVFL demonstrates a tight consensus error bound as $n$ becomes sufficiently large. As the consistency error during training decreases, the generalization error likewise decreases. This theoretical assurance underscores MvRVFL’s robust generalization performance concerning consistency.
Next, we examine the generalization error bound. According to (32), we use the weighted combination of predictions from the two views to define the prediction function for MvRVFL. Therefore, the generalization error bound for MvRVFL can be determined using the following theorem.
Theorem 2.
Given $N\in\mathbb{R}^{+}$, $\theta\in(0,1)$, and a training set $T=\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn independently and identically from probability distribution $\mathcal{D}$, where $y_{i}\in\{-1,+1\}$ and $x_{i}=(x_{i}^{A};\delta x_{i}^{B})$. Define the function class classes $\mathscr{G}=\{g|g:x\rightarrow Z^{(x)}\beta,\|\beta\|\leq N\}$ and $\hat{\mathscr{G}}=\{\hat{g}|\hat{g}:(x,y)\rightarrow yg(x),g(x)\in\mathscr{G}\}$, where $\beta=(\beta_{1};\beta_{2})$, $Z^{(x)}=\left(Z_{1}^{(x)};Z_{2}^{(x)}\right)=\left([x^{A}\hskip 5.69046pt\phi(x^{A}W^{A}+b^{A})];[x^{B}\hskip 5.69046pt\phi(x^{B}W^{B}+b^{B})]\right)$ and $g(x)=\left(Z_{1}^{(x)}\beta_{1}+\delta Z_{2}^{(x)}\beta_{2}\right)=Z^{(x)}\beta\in\mathscr{G}_{N}$. Then, with a probability of at least $1-\theta$ over $T$, every $g(x)\in\mathscr{G}$ satisfies
@@
Let’s consider a loss function $\Omega:\mathbb{R}\rightarrow[0,1]$ defined as:
$\Omega(x)=\left\{\begin{array}[]{lll}{1},&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ x<0,\vspace{3mm}\\ {1-\frac{x}{1+\delta}},&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ 0\leq x\leq 1+\delta,\\ {0},&\text{otherwise}.\end{array}\right.$ (62)
Then, we have
@@
Using Lemma 1, we have
@@
Given that $(\Omega-1)(x)$ is a Lipschitz function with a constant of $\frac{1}{1+\delta}$, passes through the origin, and is uniformly bounded, we can derive the following inequality based on Lemma 3:
$\hat{R}_{n}((\Omega-1)\circ\hat{\mathscr{G}})\leq\frac{2}{1+\delta}\hat{R}_{n}(\hat{\mathscr{G}}).$ (67)
Using the Definition 1, we obtain:
$\hat{R}_{n}(\hat{\mathscr{G}})$ $=\mathbb{E}_{\sigma}\left[\underset{\hat{g}\in\hat{\mathscr{G}}}{\sup}\left|\frac{2}{n}\sum_{i=1}^{n}\sigma_{i}\hat{g}(x_{i},y_{i})\right|\right]$
@@
Combining this with Lemma 2, we have:
$\hat{R}_{n}((\Omega-1)\circ\hat{\mathscr{G}})$ $\leq\frac{2}{1+\delta}\hat{R}_{n}(\mathscr{G})$
@@
Moreover, by combining equations Eqs. (63), (65), (A.1), and (A.1), we can obtain inequality, which demonstrates the generalization error bound of MvRVFL. ∎
@@
We define the classification error function $g(x)$ by employing the integrated decision function as specified in MvRVFL (32). By integrating the empirical Rademacher complexity of $\mathscr{G}$ with the empirical expectation of $\hat{\mathscr{G}}$, we derive a margin-based estimate of the misclassification probability. Clearly, MvRVFL provides a robust generalization error bound for classification as $n$ becomes sufficiently large. As the training error decreases, the generalization error correspondingly reduces. This theoretical result ensures that MvRVFL exhibits improved generalization performance.
A.2 Sensitivity Analysis
In this section, we perform the sensitivity analysis of several key hyperparameters of the proposed MvRVFL model. These analyses covered various factors, including hyperparameters $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ discussed in subsection A.2.1. Performance with different numbers of training samples on the AwA dataset is discussed in subsection A.2.2. Finally, we conduct the sensitivity of $\theta$ and $\rho$ discussed in subsection A.3.1.
A.2.1 Effect of the parameter $\mathcal{C}_{1}$ and $\mathcal{C}_{1}$ on the performance of the proposed MvRVFL model on AwA dataset
The performance of the proposed MvRVFL model is assessed by adjusting the values of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. This thorough analysis helps us pinpoint the configuration that enhances predictive accuracy and improves the model’s robustness against new data samples. Fig. 10 illustrates significant variations in the model’s accuracy across different values of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, underscoring the model’s sensitivity to these specific hyperparameters.
According to the findings presented in Fig. 10, optimal performance of the proposed model is observed within the $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ ranges of $10^{-4}$ to $10^{4}$. These findings indicate that both $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ significantly impact the model’s performance. Hence, it is advisable to meticulously select the hyperparameters $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ in the MvRVFL model to achieve superior generalization performance.
(a) Chimpanzee vs Giant panda
(b) Giant panda vs Leopard
(c) Persian cat vs Rat
(d) Giant panda vs Hippopotamus

### v2

@@ 5 Generalization Capability Analysis
If we multiply by $y_{i}$ on both sides of the constraints of Eq. (5), it results in
@@
$y_{i}Z_{2}^{(x_{i})}\beta_{2}-1=y_{i}\xi_{2_{i}},\hskip 5.69046pti=1,2,\ldots,n,$ (28)
Eqs. (27) and (28) are used in the theorem $5.$
To start, we define the Rademacher complexity [4] as follows:
@@
For a sample set $S=\{x_{1},\ldots,x_{n}\}$, consisting of $n$ independent samples drawn from the distribution $\mathcal{D}$, and a function class $\mathscr{G}$ defined on $S$, the empirical Rademacher complexity of $\mathscr{G}$ is expressed as:
@@
where $\sigma=(\sigma_{1},\ldots,\sigma_{n})$ represents a sequence of independent random variables, each taking values from the set $\{+1,-1\}$ with equal probability (Rademacher variables). We can express the Rademacher complexity for the function class $\mathscr{G}$ as:
@@
Lemma 5.1.
Choose $\theta$ within the interval $(0,1)$, and let $\mathscr{G}$ denote a set of functions mapping the input space $S$ to the range $[0,1]$. Suppose $\{x_{i}\}_{i=1}^{n}$ are independently sampled from a probability distribution $\mathcal{D}$. Then, with a probability of at least $1-\theta$ over random samples of size $n$, every function $g\in\mathscr{G}$ meets the specified condition.
@@
Lemma 5.3.
Suppose $\mathscr{A}$ is a Lipschitz continuous function with a Lipschitz constant $\mathcal{L}$, mapping real numbers to real numbers and satisfying $\mathscr{A}(0)=0$. The Rademacher complexity for the function class $\mathscr{A}\circ\mathscr{G}$ is given by the following expression:
@@
The discrepancy between the final outputs from the two views is represented by $g_{m}(x)=\lvert Z_{1}^{(x)}\beta_{1}-Z_{2}^{(x)}\beta_{2}\rvert$. Thus, the theoretical expectation of $g_{m}(x)$ can be derived from the following theorem.
Theorem 5.4.
Let $N\in\mathbb{R}^{+}$, $\theta\in(0,1)$, and a training set $T=\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn independently and identically from probability distribution $\mathcal{D}$, where $y_{i}\in\{-1,+1\}$ and $x_{i}=(x_{i}^{A};x_{i}^{B})$. Define the function class $\mathscr{G}_{N}=\{g|g:x\rightarrow\lvert Z^{(x)}\beta\rvert,\|\beta\|\leq N\}$ and $\hat{\mathscr{G}}_{N}=\{\hat{g}|\hat{g}:x\rightarrow Z^{(x)}\beta,\|\beta\|\leq N\}$, where $\beta=(\beta_{1};\beta_{2})$, $Z^{(x)}=\left(Z_{1}^{(x)};-Z_{2}^{(x)}\right)=\left([x^{A}\hskip 5.69046pt\phi(x^{A}W^{A}+b^{A})];-[x^{B}\hskip 5.69046pt\phi(x^{B}W^{B}+b^{B})]\right)$ and $g_{m}(x)=\lvert Z_{1}^{(x)}\beta_{1}-Z_{2}^{(x)}\beta_{2}\rvert=\lvert Z^{(x)}\beta\rvert\in\mathscr{G}_{N}$. Then, every $g_{m}(x)\in\mathscr{G}_{N}$ satisfies with a probability of at least $1-\theta$ over $T$:
@@
Let $\Omega:\mathbb{R}\to[0,1]$ be a loss function defined as follows:
$\Omega=\left\{\begin{array}[]{lll}{-\frac{x}{N\mathcal{K}_{m}}},&\text{if}\penalty 10000\ \penalty 10000\ -N\mathcal{K}_{m}\leq x<0,\vskip 8.53581pt\\ {\frac{x}{N\mathcal{K}_{m}}},&\text{if}\penalty 10000\ \penalty 10000\ 0\leq x\leq N\mathcal{K}_{m},\\ {1},&\text{otherwise}.\end{array}\right$. (39)
@@
In that case, $g_{m}(x)$ ranges from 0 to $N\mathcal{K}_{m}$, while $\hat{g}_{m}(x)$ ranges from $N\mathcal{K}_{m}$ to $N\mathcal{K}_{m}$. According to Lemma 5.1 and with $\Omega(\hat{g}_{m}(x))$ ranging from 0 to $1$, with a probability of at least $1-\theta$ over $T$, the following inequality is true:
@@
Assuming $\Omega(x)$ is a Lipschitz continuous function with a constant of $\frac{1}{N\mathcal{K}_{m}}$, passes through the origin, and is uniformly bounded, the following inequality can be established using the findings from Lemmas 5.2 and 5.3:
@@
From Eqs. (5) and (46), we can conclude the theorem. ∎
Consistency plays a crucial role in MVL, aiming to reduce the discrepancy in predictions across different perspectives or views. A lower consistency error typically results in improved generalization performance. The function $g_{m}(x)$ is introduced to quantify the consistency discrepancy between the outputs of the two views. By leveraging the empirical Rademacher complexity of the function classes $\hat{\mathscr{G}}_{m}$ and the empirical expectation of $g_{m}$, a margin-based estimate of the expected consistency gap is developed. As the sample size $n$ increases, MvRVFL achieves a tighter bound on the consensus error. As the consistency error during training decreases, the generalization error likewise decreases. This theoretical assurance underscores MvRVFL’s robust generalization performance concerning consistency.
Next, we examine the generalization error bound. According to (25), we use the weighted combination of predictions from the two views to define the prediction function for MvRVFL. Consequently, the generalization error bound for the MvRVFL model can be established as outlined in the subsequent theorem.
Theorem 5.5.
Given $N\in\mathbb{R}^{+}$, $\theta\in(0,1)$, and a training set $T=\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn independently and identically from probability distribution $\mathcal{D}$, where $y_{i}\in\{-1,+1\}$ and $x_{i}=(x_{i}^{A};\delta x_{i}^{B})$. Define the function classes $\mathscr{G}=\{g|g:x\rightarrow Z^{(x)}\beta,\|\beta\|\leq N\}$ and $\hat{\mathscr{G}}=\{\hat{g}|\hat{g}:(x,y)\rightarrow yg(x),g(x)\in\mathscr{G}\}$, where $\beta=(\beta_{1};\beta_{2})$, $Z^{(x)}=\left(Z_{1}^{(x)};Z_{2}^{(x)}\right)=\left([x^{A}\hskip 5.69046pt\phi(x^{A}W^{A}+b^{A})];[x^{B}\hskip 5.69046pt\phi(x^{B}W^{B}+b^{B})]\right)$ and $g(x)=\left(Z_{1}^{(x)}\beta_{1}+\delta Z_{2}^{(x)}\beta_{2}\right)=Z^{(x)}\beta\in\mathscr{G}_{N}$. Then, every $g(x)\in\mathscr{G}$ fulfils with a probability of at least $1-\theta$ over $T$ satisfies
@@
Let $\Omega:\mathbb{R}\to[0,1]$ be a loss function defined as follows:
$\Omega(x)=\left\{\begin{array}[]{lll}{1},&\text{if}\penalty 10000\ \penalty 10000\ x<0,\vskip 8.53581pt\\ {1-\frac{x}{1+\delta}},&\text{if}\penalty 10000\ \penalty 10000\ 0\leq x\leq 1+\delta,\\ {0},&\text{otherwise}.\end{array}\right$. (51)
Following that, we obtained
@@
Using Lemma 5.1, we have
@@
Given that $(\Omega-1)(x)$ exhibits Lipschitz continuity with a Lipschitz constant of $\frac{1}{1+\delta}$, is bounded, and passes through the origin, we can establish the following inequality using Lemma 5.3:
[…]
Next, we obtain (using the Definition 1):
[…]
@@
Combining this with Lemma 5.2, we have:
[…]
@@
Moreover, by combining equations Eqs. (52), (54), (5), and (5), we can obtain inequality, which demonstrates the generalization error bound of MvRVFL. ∎
@@
We define error function $g(x)$ by employing the integrated decision function as specified in MvRVFL (25). Integrating the empirical Rademacher complexity of $\mathscr{G}$ with the empirical mean of $\hat{\mathscr{G}}$ allows us to derive a margin-based estimation of the classification error rate. As the number of samples $n$ grows, the MvRVFL model yields a tight generalization error bound, enhancing its classification capability. A reduction in training error typically leads to a simultaneous decline in the generalization error. This theoretical result ensures that MvRVFL exhibits improved generalization performance.
6 Experiments and Results
The effectiveness of the proposed MvRVFL model is evaluated by conducting a comparative analysis with baseline models on the DNA-binding proteins dataset [36]. Furthermore, we evaluate our proposed model using publicly available AwA [^1] http://attributes.kyb.tuebingen.mpg.de and Corel5k [^2] https://wang.ist.psu.edu/docs/related/ benchmark datasets. We compare our proposed MvRVFL models with SVM2K [18], MvTSVM [70], ELM (Extreme learning machine, also known as RVFL without direct link (RVFLwoDL)) [31], RVFL [46], MVLDM [29], CDC (A simple framework for complex data clustering) [34], and MCMC (Multilevel contrastive multiview clustering with dual self-supervised learning) [5]. We denote the ELM model as ELM-Vw-$A$ and ELM-Vw-$B$ if it is trained over Vw-$A$ and Vw-$B$ of the datasets, respectively. Similar nomenclature is followed for RVFL-Vw-$A$ and RVFL-Vw-$B$.
6.1 Experimental Setup
The experiments are performed on a personal computer equipped with an Intel(R) Xeon(R) Gold 6226R processor running at 2.90 GHz. The system uses Windows 11 with 128 GB of RAM and executes Python 3.11. The dataset is randomly divided into training and testing subsets, with $30\%$ allocated for testing and $70\%$ for training. A five-fold cross-validation approach combined with grid search is utilized to fine-tune the model hyperparameters, examining the specified ranges: $\mathcal{C}_{i}=\theta=\rho=\{10^{-5},10^{-4},\ldots,10^{5}\}$ for $i=1,2.$ The range $3:20:203$ is used to choose the number of hidden nodes.
6.2 Evaluation on DNA-binding Protein Dataset
We conduct a comparison of the proposed MvRVFL model by utilizing two benchmark datasets, namely PDB186 and PDB1075. The training dataset, consisting of $1075$ protein samples, is derived from the PDB1075 [36] dataset. Within this dataset, $550$ sequences are labeled as negative (non-DBPs), while $525$ sequences are categorized as positive (DBPs). The test set, comprising $186$ protein samples, is derived from the PDB186 dataset [41]). It consists of an equal number of negative and positive sequences. The features partially dictate the upper-performance limit of the model. To evaluate the impact of various features and their combination on DBP prediction, we evaluate each individual feature utilized in the proposed MvRVFL model. The DBP sequence is represented through three distinct views: physicochemical property, evolutionary information, and amino acid composition. These views are converted into feature matrices through various extraction techniques. Specifically, the amino acid composition is processed using Multi-scale Continuous and Discontinuous (MCD) [73], the physicochemical properties are analyzed with Normalized Moreau-Broto Autocorrelation (NMBAC), and evolutionary information is transformed through methods like PSSM-based Average Blocks (PSSM-AB), PSSM-based Discrete Wavelet Transform (PSSM-DWT), and Pseudo Position-Specific Scoring Matrix (PsePSSM) [37]. Five distinct types of features extracted from the sequences are utilized, encompassing PsePSSM, PSSM-DWT, NMBAC, MCD, and PSSM-AB.
Table 3: The performance of the proposed MvRVFL-1 and MvRVFL-2 models is compared with baseline models based on classification Acc. for the DNA-binding protein datasets.
Dataset Dataset SVM2K [18] MvTSVM [70] ELM-Vw-AA [31] ELM-Vw-$B$ [31] RVFL-Vw-AA [46] RVFL-Vw-$B$ [46] MVLDM [29] MvRVFL-1 MvRVFL-2
index name (Acc., Seny.) (Acc., Seny.) (Acc., Seny.) (Acc., Seny.) (Acc., Seny.) (Acc., Seny.) (Acc., Seny.) (Acc., Seny.) (Acc., Seny.)
Average Acc. $68.45$ $69.04$ $68.06$ $68.12$ $73.49$ $70.06$ $63.82$ 76.32 $74$
Seny. denote the Sensitivity.
represents the proposed models.
Table 3 provides the results (in terms of Accuracy and Sensitivity) among the baseline and proposed MvRVFL models for DBP prediction. The MvRVFL-1 and MvRVFL-2 models ranked first and second, respectively, with average accuracies of 76.32% and 74%. In contrast, the average Acc. of the baseline SVM2K, MvTSVM, ELM-Vw-$A$, ELM-Vw-$B$, RVFL-Vw-AA, RVFL-Vw-$B$, and MVLDM models are $68.45\%$, $69.04\%$, $68.06\%$, $68.12\%$, $73.49\%$, $70.06\%$ and $63.82\%$, respectively. Compared to the third-top model, RVFL-Vw-$A$, the proposed models (MvRVFL-1 and MvRVFL-2) exhibit average Acc. of approximately $2.83\%$ and $0.51\%$ higher, respectively. For the MCD & PsePSSM, NMBAC & PsePSSM, PsePSSM & PSSM-AB and PsePSSM & PSSM-DWT cases, the proposed MvRVFL-1 model archives the Acc. of $11\%, 80.80$$80\%$, $75.65\%$, and $74.41\%$, respectively, emerging as the top performers. The results suggest the significance of PsePSSM as an important feature for predicting DBPs. Thus, the proposed MvRVFL consistently demonstrates exceptional performance, achieving high Acc. across diverse scenarios, solidifying its distinction among other models. From Table 3, we can see that our proposed MvRVFL-1 model has Seny. $89.25$ which is the second highest among all the datasets. The MvRVFL-1 model is highly effective at correctly detecting true positive cases in the MCD & NMBAC dataset, thus demonstrating its superior performance in identifying relevant instances compared to other models. Also, the Seny. for the MCD & PSSM-DWT dataset reaches a peak value of $91.40$, indicating that our proposed model excels in predicting MCD features, achieving the best performance among all compared models. Furthermore, the MvRVFL-2 model employs a majority voting mechanism, combining predictions from multiple classifiers to enhance its efficiency and robustness. This approach leads to more reliable and accurate predictions, underscoring the superior performance of the MvRVFL-2 model in predicting MCD features compared to baseline models.
Figure 4: ROC curves comparing the performance of the proposed MvRVFL model with various DNA-binding protein prediction models.
We compare our proposed MvRVFL model with existing DNA binding protein prediction models, including MvLSSVM via HSIC [81], HKAM-MKM [80], MV-H-RKM [25], MLapSVM-LBS [59], and LapLKA-RKM [49]. Table 4 demonstrates that all five features (MCD, NMBAC, PSSM-DWT, PsePSSM, and PSSM-AB) are beneficial for predicting DBPs using MvRVFL. Among these, MCD_vs_PsePSSM achieves the best result (Acc. = $80.11\%$). This suggests that PsePSSM is a crucial feature for predicting DBPs. As illustrated in Table 4, the proposed MvRVFL-1 and MvRVFL-2 models achieve average accuracies of 76.32% and 74%, respectively, both outperforming the baseline models. The proposed MvRVFL-1 and MvRVFL-2 exhibit exceptional generalization performance, marked by consistently higher Acc., indicating a high level of confidence in their learning process. It is found that the MvRVFL model is comparable to the baseline models in most cases. By incorporating the coupling term, MvRVFL can integrate information from both views. This approach allows for a larger error variable in one view if it is compensated by the other view by minimizing the product of the error variables. The MvRVFL-1 model demonstrates strong performance in all these areas, showing its robustness and reliability in comparison with existing models.
Figure 4 displays the ROC curve, highlighting the enhanced performance of the proposed MvRVFL model over baseline methods on DNA-binding protein datasets. The curve offers a thorough assessment of the model’s classification performance by plotting sensitivity (true positive rate) versus the false positive rate across different decision thresholds. The MvRVFL model achieves a significantly higher area under the ROC curve (AUC), indicating a better balance between sensitivity and specificity. An elevated AUC indicates that the MvRVFL model possesses a stronger capability to differentiate positive cases from negative ones, resulting in improved prediction accuracy. The increase in AUC demonstrates the model’s improved ability to minimize false negatives while accurately identifying true positives, leading to more reliable detection.
These results demonstrate the MvRVFL model’s superiority over baseline models in classification tests, demonstrating its superiority and effectiveness. The importance of PsePSSM as a feature is evident, as it plays a pivotal role in improving the model’s overall accuracy and reliability in identifying true positives and reducing false negatives.
Figure 5: Ablation study of the coupling term for MvRVFL: The x-axis represents the datasets indexed in Table 4, and the y-axis represents the Acc. ($\%$).
6.2.3 Ablation study on DNA binding datasets
This result underscores the importance of the coupling term $\rho$ in the model’s design. It enables the MvRVFL model to seamlessly combine information from both the image and caption views. By reducing the product of the error terms from both views, the model can tolerate a higher error in one view, as long as it is compensated by a lower error in the other view. This balancing approach improves the model’s overall robustness and reliability.
Dataset Dataset MvLSSVM via HSIC [81] HKAM-MKM [80] MV-H-RKM [25] MLapSVM-LBS [59] LapLKA-RKM [49] MvRVFL-1 MvRVFL-2
index name (Acc., Seny.) (Acc., Seny.) (Acc., Seny.) (Acc., Seny.) (Acc., Seny.) (Acc., Seny.) (Acc., Seny.)
Average Acc. $73.67$ $75.07$ $73.45$ $70.92$ $71.34$ 76.32 $74$
Seny. denote the Sensitivity.
represents the proposed models.
6.2.4 Sensitivity of $\mathcal{C}_{1}$ and $\rho$ combinations
In this subsection, we evaluate the impact of the regularization parameters $\mathcal{C}_{1}$ and $\rho$ of MvRVFL. In Fig. 6, the values of $\mathcal{C}_{1}$ and $\rho$ are varied from $10^{-5}$ to $10^{5}$, keeping the other parameters is fixed at their optimal values and the corresponding Acc. values are recorded. The performance of MvRVFL is depicted under different parameter settings $\mathcal{C}_{1}$ and $\rho$. The parameter $\rho$ serves as a coupling term designed to minimize the product of error variables between Vw-$A$ and Vw-$B$. When $\mathcal{C}_{1}$ lies between $10^{-1}$ and $10^{5}$ with the same value of $\rho$, there is a corresponding improvement observed in the Acc. values. These findings suggest that when evaluating parameters $\mathcal{C}_{1}$ and $\rho$, the model’s performance is primarily influenced by $\mathcal{C}_{1}$ rather than $\rho$. Therefore, precise adjustment of the hyperparameters in the proposed MvRVFL model is essential for achieving optimal generalization performance.
6.2.5 Influence of the number of hidden nodes $h_{l}$
To thoroughly understand the robustness of the proposed MvRVFL model, it is essential to examine how it responds to variations in the number of hidden nodes $h_{l}$. The influence of the hyperparameter $h_{l}$ is depicted in Fig. 7. Fig. 7(a) and 7(c) show that the performance peaks at $h_{l}=23$ and then gradually declines as $h_{l}$ increases further. Therefore, to achieve the best performance from the MvRVFL, we recommend using $h_{l}=23$. The performance peaks at $h_{l}=23$ and $h_{l}=83$ and then gradually declines as $h_{l}$ increases further depicted in Fig. 7(d). From Fig. 7(b), the performance consistently improves with an increase in the number of hidden nodes until reaching a plateau. Optimal performance is typically achieved with higher values of $h_{l}$. We recommend fine-tuning the hyperparameters to attain the best performance from the proposed models for specific tasks.